Sir Donald Bradman had scored 6996 runs in 69 completed test innings. He had one more innings to play. How many runs did he need to score in his final innings in order that his test batting average was greater than or equal to 100?
(batting = Total runs scored divided by total innings completed)

Answer by dkppathak(439)

total score was 6996 in 69 test
new test play will have total test =69+1=70
to have average score greater than or equal to 100
total score should be 100x70 = 7000 score
run required in final match will be 7000-6996 =4
therefore to get average of 100 runs
he should make the run at least 4 or more
ANSWER 4 run or more